# Spatial structure facilitates evolutionary rescue by cost-free drug resistance

### v1

@@ 2.2 Simulation runs and analyses
In general, we run simulations as follows. For each drug addition time $t_{add}$ considered, we run multiple simulation replicates. In each of them, we initialize the system with $0.1\times K$ sensitive individuals in each deme (or $0.1\times DK$ in the well-mixed population), let the system evolve until $t_{add}$ and save its state. Then we add the biostatic drug, and we let the system evolve until either extinction of the population or full system colonization by resistant individuals. If resistance fixes before drug addition, we save the state at this point and stop the simulation. When stopping the simulation, in all cases described above, we save the stopping time. Because of stochastic fluctuations around the mean, we consider that mutants have colonized one deme (resp. the whole population in the well-mixed case) if the population size is greater than $0.9\times K$ (resp. $0.9\times DK$).
@@
In Figure 2, we calculate the survival probability as the fraction of simulation replicates where bacteria survive the application of the drug, in the sense that one deme comprises at least $0.9\times K$ mutants. Note that this allows us to exclude cases in which mutants quickly go stochastically extinct after drug addition. In the structured populations, survival is intended to be in at least one of the demes.
@@
In Figure 3A-B, for each time and each structure, we perform simulation replicates in the absence of drug. For each of these replicates, we count the number of mutants in the population. We then calculate the average and variance of the total number of mutants across the different replicates.
@@
In Figure S4, for each time and each structure, we perform simulation replicates in the absence of drug. For each replicate we choose one deme (deme 1 in practice in Figure S4, but the result is similar for other demes given the symmetry of the clique). We calculate the average and variance of the number of mutants in that deme across stochastic realizations.
@@
In Figure 3C, we analyze the composition of the population before the addition of drug. In the state of the system saved before environmental change, we focus on the number of mutants, and partition the outcomes of our stochastic simulations into the four categories described in the main text (“Only R”, “Big R pop.”, “Small R pop.”, “No R”). We report the fraction of replicates falling in each of these categories.
@@
In Figure 4, for each time and a structure, we perform multiple replicates. For each of them, we calculate the number of demes where the mutant type has fixed before drug addition. We report a histogram of these number of demes across replicates.
@@
In Figure 5A, we perform simulations to assess the time until colonization of next deme. These simulation are set up slightly differently than others, to focus on the colonization process. We initialize one of the demes with $N=0.9\times K$ mutants, all others with $N=0.9\times K$ wild-types, mimicking fixation of mutants in one deme before drug addition. We add the biostatic drug at $t=0$ in this simulation, thus modeling the case where one deme has fixed resistance before drug addition. We then let the system evolve until overall colonization of the structured population by resistant mutants. For each $k\geq 1$, we save the times of successful colonization of the $(k+1)^{th}$ deme, given that $k$ demes were already colonized by mutants. We define successful colonization of a deme as reaching a number of mutants $0.9\times K$. We then obtain the values of $\tau_{c\,mig}(k)$ as the differences between the successive times we recorded.
@@
In Figure 5B, we report the time until overall colonization of the structured population by resistant mutants. When saving the state of the system after overall mutant colonization (resp. extinction), we also record the time it took for overall colonization (resp. extinction). Thus, we can calculate the mean of the overall colonization time for each $t_{add}$ and each structure. Note that the fixation of resistance may happen before drug addition (this happens in particular for addition times longer than $\langle t_{af\,S}\rangle$). In this case, we consider that colonization happens when resistant individuals reach fixation in all demes and there are at least $0.9\times K$ resistant individuals.
@@ 3.1 Impact of a cost or of a benefit of resistance
In the main text, we considered the case where the resistant mutant is neutral in the absence of drug. However, antibiotic resistance often carries a cost in the absence of drug [Andersson99, Andersson10]. Besides, if antibiotic is already present at relatively low doses in the environment before the addition of drug, it can induce a benefit of resistance before drug is added. Here, we briefly discuss these two cases.
Cost of resistance.
A cost of resistance biases the competition between mutants and wildtypes, making it difficult for mutants to fix once they have appeared. We show in Figure S1 the survival probability of a spatially structured population after drug treatment, in the case where the fitness of the mutant is $f_{R}=0.9$ before drug addition (while $f_{S}=1$). We consider both a clique (i.e. the minimal model of spatial structure considered in most of the present work), and a star with different migration asymmetries $\alpha$. These results can be compared to those in the cost-free resistance case, shown in Figure 2 for the clique, and in Figure S5 for the star. We consider a similar range of drug addition time $t_{add}$ as in those figures. In Figure S1, we do not observe a statistically significant impact of population structure on the survival probability of the population.
@@
With a cost of resistance $\delta=(f_{S}-f_{R})/f_{S}>0$ (giving $\delta=0.1$ in Figure S1) satisfying $\delta\ll 1$ and $N\delta\gg 1$ (satisfied in Figure S1), the probability of fixation of one mutant in a deme of size $N$ is approximately $\delta e^{-N\delta}$ in the diffusion approximation. It is thus exponentially suppressed, and the time to appearance of a successful mutant would be much longer than in the neutral case, on average of $\langle t_{af}\rangle\approx e^{N\delta}/(N\mu g\delta)$, giving $\langle t_{af}\rangle\approx 2.4\times 10^{13}$ with the parameter values of Figure S1. Therefore, fixation events can be neglected in the range of $t_{add}$ considered in Figure S1. Thus, we can estimate the survival probability from the probability of presence of mutant lineages destined for extinction, introduced in Equation (S16). This analytical prediction agrees with our simulation result, see Figure S1.
Still for resistant mutants with cost $\delta$ satisfying $\delta\ll 1$ and $N\delta\gg 1$, the average fixation time $\langle t_{af}\rangle\approx e^{N\delta}/(N\mu g\delta)$ of a successful mutant increases as $N$ increases. Thus, the mean fixation time of a successful mutant is substantially shorter in an average deme than in a well-mixed population. Therefore, if times comparable to the mean fixation time of a successful mutant in a deme (or larger) were considered, we predict that spatial structure would favor population survival even more than in the neutral case. While this is a potentially interesting effect, we do not consider it further here because of the extremely long times involved.
@@
Figure S1: Survival probability of a structured bacterial population with a cost of resistance. The survival probability of the population to the addition of biostatic drug is shown versus the drug addition time. We consider a clique and a star with different values of migration asymmetry $\alpha$. The case of the well-mixed population with the same total size (“WMP”) is shown for reference. The migration rate is $\gamma=10^{-6}$ in the clique, while in the star we follow the convention in Section 4 to compare with a clique of a given migration rate $\gamma$, leading to $\gamma_{O}=(D-1)\gamma$ and $\gamma_{I}=\alpha\gamma_{O}$. Simulation results are obtained from $10^{3}$ replicates, with error bars representing 95% confidence intervals. Red horizontal dashed line (“Theory”): analytical prediction from Equation (S16). Parameter values: $K=200$, $D=5$, $f_{S}=1$ without drug, $f_{S}=0$ with drug, $f_{R}=0.9$ with and without drug, $g=0.1$, $\mu=10^{-5}$.
@@
Let us now consider the case where the resistant mutant has a selective advantage $s=(f_{R}-f_{S})/f_{S}>0$ in the absence of drug. Assuming $Ns\gg 1$, the probability of fixation of one mutant in a deme of size $N$ is approximately $s$ in the diffusion approximation. As this is independent of $N$, the same holds for a well-mixed population. Thus, in this case, the average time to appearance of a successful mutant is the same in a deme and in a well-mixed population. This entails that the effect of spatial structure we evidenced for neutral and effectively neutral resistant mutants does not extend to beneficial ones.
@@ 3.2 Stochastic extinction
where $\mathbf{R}$ is the transition rate matrix describing the population dynamics, see Ref. [marrec_resist_2020]. In particular, for a population with a carrying capacity $K=10^{2}$, the probability of rapid initial extinction starting from 10 individuals or more is negligible (lower than $10^{-9}$) [marrec_resist_2020]. This is why, in our categorization of results used in Figure 3, we called “Large R population” the case where the number of mutants is at least 10: the possibility of stochastic extinction can be neglected in this case.
3.3 Growth of mutant population in structured and well-mixed populations
Starting from an inoculum of sensitive bacteria, resistant mutants may appear and fix in the absence of drug. In the case of a structured population with small migration rates, they locally fix deme after deme. This results in a substantial difference in the growth pattern of the number of resistant individuals between a well-mixed and a structured population. We present one specific trajectory of the number of sensitive and resistant individuals in a well-mixed population in Figure S2A and one in a structured population in Figure S2B. In the structured population, a locally successful mutant appears on average earlier than a successful one in the well-mixed population. However, it then takes more time for resistance to fix in the whole population, as this involves migration of resistant mutants from deme to deme (or independent appearance of other mutants if migration rates are extremely small). Accordingly, Figure S2A features a rapid but late growth of mutant fraction in the well-mixed population. Meanwhile, Figure S2B, we observe that the growth of mutant fraction in the structured population features $D-1$ intermediate plateaus, as fixation occurs deme after deme.
Despite these differences in the individual trajectories of the number of mutants, their average over many replicates is the same whatever the migration rate, see Figure 3A. However, these differences in trajectories are at the root of those observed in the variance across replicates shown in Figure 3B. The well-mixed population features more variance in mutant numbers across replicates than structured populations with small migration rates, because of the variability of the appearance time of a successful mutant. More precisely, the appearance time of a successful mutant in the fastest deme of a structured population has smaller variance, as it is exponentially distributed with a larger rate, see Equation S9, compared to Equation S5 which holds for a well-mixed population.
Figure S2: Dynamics of population in a well-mixed and a structured population. We present one trajectory from a single simulation realization in the absence of drug for a well-mixed population in Panel A, and one for a structured population with same total size in Panel B. In both cases, we show the number of sensitive (S) and resistant (R) individuals versus time. Parameter values: $K=200$, $D=5$, $\mu=10^{-5}$, $f_{S}=f_{R}=1$, $g=0.1$, $\gamma=10^{-7}$.
3.4 Population composition as a function of time for different migration rates
In Figure 3C, we analyzed the composition of the population at the time featuring the largest inter-replicate variance for the well-mixed population, namely $t=5\times 10^{5}$, in spatially structured populations with different migration rates. In Figure S3, we report how this composition evolves in time, in the absence of drug, for spatially structured populations with different migration rates. As in Figure 3C, we focus on the number of mutants in each deme to categorize the population composition.
Figure S3: Population composition versus time for different population structures. Population composition in the absence of drug in the well-mixed population (“WMP”, Panel A), in the fully subdivided population (“SP”, Panel B), and in the clique-structured population with same total size for three values of the migration rate $\gamma$ (Panels C-E). The four categories of population composition reported here are based on the number of mutants per deme. They are defined in the main text and used in Figure 3C. The x-axis tick corresponding to $t=5\times 10^{5}$, considered in Figure 3C, is highlighted in purple. Data is obtained from $10^{4}$ simulation replicates in each case. Parameter values (in all panels): $K=100$, $D=10$, $f_{S}=f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
Recall that here, the population is composed of wild-type sensitive bacteria and of neutral resistant mutants that arise with probability $\mu$ upon division of wild-types. Figure S3 shows that when the per capita migration rate $\gamma$ is decreased, there is a longer phase where the population has an intermediate composition, with some demes harboring a substantial number of resistant mutants (“Big R pop.” category). This is consistent with our above discussion in Section 3.3 and with our results in Figure S2. The fixation of mutants occurs gradually, deme by deme, in structured populations with small migration rates. Furthermore, substantial mutant numbers exist earlier in these structured populations. We further observe in Figure S3 that a per capita migration rate $\gamma=10^{-3}$ results in an evolution of population composition akin to that of a well-mixed population, a per capita migration rate $\gamma=10^{-7}$ yields an evolution of population composition similar to that of a fully subdivided population with $\gamma=0$. Structures with intermediate values of $\gamma$ feature an intermediate evolution of population composition.

### v2

@@ 2.2 Simulation runs and analyses
In general, we run simulations as follows. For each drug addition time $t_{\textrm{add}}$ considered, we run multiple simulation replicates. In each of them, we initialize the system with $0.1\times K$ sensitive individuals in each deme (or $0.1\times DK$ in the well-mixed population), let the system evolve until $t_{\textrm{add}}$ and save its state. Then we add the biostatic drug, and we let the system evolve until either extinction of the population or full system colonization by resistant individuals. If resistance fixes before drug addition, we save the state at this point and stop the simulation. When stopping the simulation, in all cases described above, we save the stopping time. Because of stochastic fluctuations around the mean, we consider that mutants have colonized one deme (resp. the whole population in the well-mixed case) if the population size is greater than $0.9\times K$ (resp. $0.9\times DK$).
@@
In Fig 2, we calculate the survival probability as the fraction of simulation replicates where bacteria survive the application of the drug, in the sense that one deme comprises at least $0.9\times K$ mutants. Note that this allows us to exclude cases in which mutants quickly go stochastically extinct after drug addition. In the structured populations, survival is intended to be in at least one of the demes.
@@
In Fig 3A-B, for each time and each structure, we perform simulation replicates in the absence of drug. For each of these replicates, we count the number of mutants in the population. We then calculate the average and variance of the total number of mutants across the different replicates.
@@
In Fig S4, for each time and each structure, we perform simulation replicates in the absence of drug. For each replicate we choose one deme (deme 1 in practice in Fig S4, but the result is similar for other demes given the symmetry of the clique). We calculate the average and variance of the number of mutants in that deme across stochastic realizations.
@@
In Fig 3C, we analyze the composition of the population before the addition of drug. In the state of the system saved before environmental change, we focus on the number of mutants, and partition the outcomes of our stochastic simulations into the four categories described in the main text (“Only R”, “Big R pop.”, “Small R pop.”, “No R”). We report the fraction of replicates falling in each of these categories.
@@
In Fig 4, for each time and a structure, we perform multiple replicates. For each of them, we calculate the number of demes where the mutant type has fixed before drug addition. We report a histogram of these number of demes across replicates.
@@
In Fig 5A, we perform simulations to assess the time until colonization of next deme. These simulation are set up slightly differently than others, to focus on the colonization process. We initialize one of the demes with $N=0.9\times K$ mutants, all others with $N=0.9\times K$ wild-types, mimicking fixation of mutants in one deme before drug addition. We add the biostatic drug at $t=0$ in this simulation, thus modeling the case where one deme has fixed resistance before drug addition. We then let the system evolve until overall colonization of the structured population by resistant mutants. For each $k\geq 1$, we save the times of successful colonization of the $(k+1)^{th}$ deme, given that $k$ demes were already colonized by mutants. We define successful colonization of a deme as reaching a number of mutants $0.9\times K$. We then obtain the values of $\langle t_{\textrm{c mig}}(k)\rangle$ as the differences between the successive times we recorded.
@@
In Fig 5B, we report the time until overall colonization of the structured population by resistant mutants. When saving the state of the system after overall mutant colonization (resp. extinction), we also record the time it took for overall colonization (resp. extinction). Thus, we can calculate the mean of the overall colonization time for each $t_{\textrm{add}}$ and each structure. Note that the fixation of resistance may happen before drug addition (this happens in particular for addition times longer than $\langle t_{af\,S}\rangle$). In this case, we consider that colonization happens when resistant individuals reach fixation in all demes and there are at least $0.9\times K$ resistant individuals.
@@ 3 Clique population structure with sensitive inoculum
3.1 Clique population structure with different numbers of demes
To assess the impact of the degree of subdivision of a population on rescue by resistance, we extend the study of the survival probability with neutral R mutants in Fig 2A to different numbers of demes. Fig S1 shows the impact of varying the degree of subdivision of a population of fixed total size $DK$, but composed of a different number of demes ($D=5,10,20$). We observe that higher subdivision yields higher probabilities of survival.
Figure S1: Survival probability of a bacterial population with neutral mutants for varying degree of subdivision at a fixed total population size. Survival probability as a function of the treatment addition time $t_{\textrm{add}}$. Three levels of subdivision of a population with total carrying capacity $DK=1000$, into 5, 10, and 20 demes, are shown for two values of $\gamma$. Results for the well-mixed population (“WMP”) are shown for reference. In all cases, we determined the survival probability from the fraction of $10^{4}$ simulations where the population survived after drug application. The vertical dash-dotted lines denote the average time of appearance of a locally successful mutant in the fastest deme of each structure (Eq 2). The vertical dotted line represents the average appearance time of a successful mutant in a well-mixed population (Eq 1). Solid lines for the fully subdivided ($\gamma=0$) and well-mixed populations are analytical predictions from Eq 4. Thin lines connecting data points are guides for the eye. Parameter values: $f_{S}=1$ without drug, $f_{S}=0$ with drug, $f_{R}=1$ (no cost of resistance), $g=0.1$, $\mu=10^{-5}$.
@@ 3.2 Impact of a cost or of a benefit of resistance
In Fig 2, we considered the case where the resistant mutant is neutral in the absence of drug and the case where it carries a moderate cost of resistance $\delta=1/K$. Here, we discuss different regimes of cost of resistance, and present simulation results for a higher cost of resistance. Besides, if antibiotic is already present at relatively low doses in the environment before the addition of drug, it can induce a benefit of resistance before drug is added. We briefly discuss this case at the end of this section.
Different regimes of cost.
Let us consider different regimes of cost for R mutants.
As discussed in the main text, for neutral resistant mutants ($\delta=0$), $p_{\textrm{fix}}=1/N^{*}$, so Eq 1 gives $\langle t_{af\,W}(N^{*})\rangle=1/(\mu g)$. Importantly, this result does not depend on population size $N^{*}$. Therefore, it holds both for each deme in a structured population and for a well-mixed population of steady-state size $DN^{*}$: $\langle t_{af\,W}(DN^{*})\rangle=1/(\mu g)$. Meanwhile, Eq 2 gives $\langle t_{af\,F}(N^{*},D)\rangle=1/(D\mu g)=\langle t_{af\,W}(N^{*})\rangle/D$. Therefore, a locally successful mutant appears in a structured population $D$ times faster than a successful mutant in a well-mixed population with same total steady-state size $DN^{*}$.
For mutants with a substantial cost of resistance $\delta\gg 1/K$, $p_{\textrm{fix}}=\delta e^{-N^{*}\delta}$ (see Eq S13), and thus $\langle t_{af\,W}(N^{*})\rangle=e^{N^{*}\delta}/(N^{*}\mu g\delta)$, which is exponentially longer than in the neutral case and satisfies $\langle t_{af\,W}(N^{*})\rangle\ll\langle t_{af\,W}(DN^{*})\rangle$. Therefore, $\langle t_{af\,F}(N^{*},D)\rangle=\langle t_{af\,W}(N^{*})\rangle/D\ll\langle t_{af\,W}(DN^{*})\rangle$.
Finally, mutants with intermediate cost satisfying $1/(DK)\ll\delta\ll 1/K$ are effectively neutral in demes of size $N^{*}$, leading to $\langle t_{af\,F}(N^{*},D)\rangle=1/(D\mu g)$ as in the neutral case (see Eq S12). However, they are substantially deleterious in a well-mixed population of size $DN^{*}$, leading to $\langle t_{af\,W}(DN^{*})\rangle=e^{DN^{*}\delta}/(DN^{*}\mu g\delta)$, which is exponentially larger than $\langle t_{af\,F}(N^{*},D)\rangle$.
Strong cost of resistance.
@@
With a strong cost of resistance $\delta$ satisfying $\delta\ll 1$ and $N^{*}\delta\gg 1$ (satisfied in Fig S2), the probability of fixation of one mutant in a deme of size $N^{*}$ is approximately $\delta e^{-N^{*}\delta}$ (cf. Eq S13). It is thus exponentially suppressed, and the time to appearance of a successful mutant in a well-mixed population of size $N^{*}$ is exponentially longer than in the neutral case. With the parameter values of Fig S2, the appearance time of a successful mutant in a well-mixed population of size $DN^{*}$ is $\langle t_{af\,W}(DN^{*})\rangle\approx 1.3\times 10^{43}$, while the appearance time of the fastest locally successful mutant in a population of $D$ demes of size $N^{*}$ each is $\langle t_{af\,F}(N^{*},D)\rangle\approx 1.3\times 10^{8}$. Therefore, fixation events can be neglected in the range of $t_{\textrm{add}}$ considered previously for neutral mutants and mutants with moderate cost (see Fig 2 for the clique, and in Fig S6 for the star). Thus, in this range of $t_{\textrm{add}}$, we can estimate the survival probability from the probability of presence of mutant lineages destined for extinction, introduced in Eq S4. This analytical prediction agrees with our simulation results, see Fig S2. We do not observe a statistically significant impact of population structure on the survival probability of the population, in this range of $t_{\textrm{add}}$. Recall however that if times comparable to the mean fixation time of a successful mutant in a deme (or larger) were considered, we predict that spatial structure would favor population survival for strong costs of resistance. We do not consider this further because of the extremely long times involved.
@@
Figure S2: Survival probability of a structured bacterial population with a cost of resistance. The survival probability of the population to the addition of biostatic drug is shown versus the drug addition time. We consider a clique and a star with different values of migration asymmetry $\alpha$. The case of the well-mixed population with the same total size (“WMP”) is shown for reference. The migration rate is $\gamma=10^{-6}$ in the clique, while in the star we follow the convention in Section 4 to compare with a clique of a given migration rate $\gamma$, leading to $\gamma_{O}=(D-1)\gamma$ and $\gamma_{I}=\alpha\gamma_{O}$. Simulation results are obtained from $10^{3}$ replicates, with error bars representing 95% confidence intervals. Red horizontal dashed line (“Theory”): analytical prediction from Eq S4. Parameter values: $K=200$, $D=5$, $f_{S}=1$ without drug, $f_{S}=0$ with drug, $f_{R}=0.9$ with and without drug, $g=0.1$, $\mu=10^{-5}$.
@@
Let us now consider the case where the resistant mutant has a fitness $f_{R}=1+s$ with a selective advantage $s>0$ in the absence of drug. Assuming $s\ll 1$ but $N^{*}s\gg 1$, the probability of fixation of one mutant in a deme of size $N^{*}$ is approximately $s$ in the Moran model [ewens_mathematical_2004]. As this is independent of $N^{*}$, the same holds for a well-mixed population of size $DN^{*}$. Thus, in this case, the average time to appearance of a successful mutant is the same in a deme and in a well-mixed population. This entails that the effect of spatial structure we evidenced for neutral and deleterious resistant mutants does not extend to beneficial ones.
3.3 Growth of mutant number in structured and well-mixed populations
Starting from an inoculum of sensitive bacteria, resistant mutants may appear and fix in the absence of drug. In the case of a structured population with small migration rates, they locally fix deme after deme. This results in a substantial difference in the growth pattern of the number of resistant individuals between a well-mixed and a structured population. We present one specific trajectory of the number of sensitive and neutral resistant individuals in a well-mixed population in Fig S3A and one in a structured population in Fig S3B. In the structured population, a locally successful mutant appears on average earlier than a successful one in the well-mixed population. However, it then takes more time for resistance to fix in the whole population, as this involves either independent appearance of other mutants or migration of resistant mutants from deme to deme. Accordingly, Fig S3A features a rapid but late growth of mutant fraction in the well-mixed population. Meanwhile, Fig S3B, we observe that the growth of mutant fraction in the structured population features $D-1$ intermediate plateaus, as fixation occurs deme after deme. Note that the sequential fixation pattern shown in Fig S3 is also observed for a successful mutant that carries a cost of resistance, but the involved timescales differ (see Supplementary Appendix Section 3.2).
Despite these differences in the individual trajectories of the number of mutants, their average over many replicates is the same whatever the migration rate for neutral mutants, see Fig 3A. However, these differences in trajectories are at the root of those observed in the variance across replicates shown in Fig 3B. The well-mixed population features more variance in mutant numbers across replicates than structured populations with small migration rates, because of the variability of the appearance time of a successful mutant.
Figure S3: Dynamics of population in a well-mixed and a structured population. We present one trajectory from a single simulation realization in the absence of drug for a well-mixed population in Panel A, and one for a structured population with same total size in Panel B. In both cases, we show the number of sensitive (S) and resistant (R) individuals versus time. Parameter values: $K=200$, $D=5$, $\mu=10^{-5}$, $f_{S}=f_{R}=1$ (neutral mutants), $g=0.1$, $\gamma=10^{-7}$.
@@ 3.4 Stochastic extinction
where $\mathbf{R}$ is the transition rate matrix describing the population dynamics, see Ref. [marrec_resist_2020]. In particular, for a population with a carrying capacity $K=10^{2}$ composed of individuals with fitness $f_{R}=1$, the probability of rapid initial extinction starting from 10 individuals or more is negligible (lower than $10^{-9}$) [marrec_resist_2020]. (Note that this extinction probability can also be calculated in the branching process approximation of Section 1.1, giving $(g/f_{R})^{10}=10^{-10}$ here). This is why, in our categorization of results used in Fig 3, we called “Large R population” the case where the number of mutants is at least 10: the possibility of stochastic extinction can be neglected in this case.